1. The Metro Electric Company produces and distributes electricity to residential customers in New Brunswick.  The demand for electricity is given by P = 0.04 – 0.01Q The resulting marginal revenue function is: MR = 0.04 – 0.02Q Metro’s marginal cost function is: MC = 0.005 + 0.0075Q
  2. What is the gain in surplus from regulating the Metro’s price and how much is the monopoly worth to Metro (i.e. how much will they spend lobbying the government not to regulate their price)?
Answer #1

Monopoly produces upto the point where MC = MR, ie

0.04 - 0.02Q = 0.005 + 0.0075Q

0.035 = 0.0275Q

Q = 1.27

P = 0.04 - 0.01*1.27

P = 0.0273